6)                 If f(x) takes the values -21, 15, 12 and 3 respectively when x assumes the values -1, 1, 2 and 3, find the polynomial which approximates f(x).



1
Expert's answer
2021-07-18T10:22:25-0400

Solution.

Each of the points (-1,-21), (1,15),(2,12),(3,3​) satisfies the equation "p=ax^3+bx^2+cx+d"  for some unknown "a,b,c,d."  Substitute each point in the equation and make a matrix equation. Let be "A=\begin{pmatrix} -1& 1 &-1&1 \\ 1 & 1 & 1&1\\ 8 & 4 &2& 1\\27&9&3&1 \end{pmatrix}," "X=\begin{pmatrix} a \\ b\\ c\\d \end{pmatrix} \text{and }B=\begin{pmatrix} -21 \\ 15\\ 12\\3 \end{pmatrix}." We will have eqution


"AX=B, \text{or}
\newline\begin{pmatrix} -1& 1 &-1&1 \\ 1 & 1 & 1&1\\ 8 & 4 &2& 1\\27&9&3&1 \end{pmatrix}\cdot \begin{pmatrix} a \\ b\\ c\\d \end{pmatrix} =\begin{pmatrix} -21 \\ 15\\ 12\\3 \end{pmatrix}."

Solve matrix equation: "X = A^{-1} \u00b7 B."

"\det A=\begin{vmatrix} -1& 1 &-1&1 \\ 1 & 1 & 1&1\\ 8 & 4 &2& 1\\27&9&3&1 \end{vmatrix}=48."

"A^{-1}=\begin{pmatrix}
 -\frac{1}{6} & -\frac{1}{2} &\frac{1}{6}&0\\
 \frac{3}{4} & 2&-1&\frac{1}{4}\\
-\frac{7}{12}&-\frac{1}{2}&\frac{5}{6}&-\frac{1}{4}\\
-\frac{1}{2}&-3&-2&-\frac{1}{2}
\end{pmatrix}."

"X=\begin{pmatrix} 1 \\ -9\\ 17\\6 \end{pmatrix}, \text{or } a=1, b=-9, c=17,d=6."

So, "p=x^3-9x^2+17x+6."  It is the polynomial function.